iii) The p.d.e.
auxx + 2b uxy + c uyy = 0
where a, b ,c are constants is irreducible when
b 2− ac =0 .
1
Expert's answer
2020-03-06T12:06:57-0500

Let's rewrite the equation in operational form

"F(D,D')=aD^2+2bDD'+cD'^2=0"

An equation is irreductable if it can't be written as a product of linear factors, hence

"F(D;D')=(a_1D+b_1D'+c_1)(a_2D+b_2D'+c_2)=a_1a_2D^2+"

"+(a_1b_2+a_2b_1)DD'+b_1b_2D'^2+(a_1c_2+c_1a_2)D+(b_1c_2+c_1b_2)D'+c_1c_2"

Using the method of undetermined coefficients, we assume

"a_1a_2=a; b_1b_2=c; a_1b_2+a_2b_1=2b;"

"b_1c_2+c_1b_2=a_1c_2+c_1a_2=c_1c_2=0"

Hence "c_1=c_2=0"

If "b^2-ac=0," "a_1=a_2=\alpha; b_1=b_2=\beta" ,therefore "F(D,D')=(\alpha D+\beta D')^2" which is irreducible